# A Unified Model for Multi-Task Drone Routing in Post-Disaster Road Assessment

### v1

@@ 5.1 Model architecture
Figure 2: Overall Model Architecture for Unified Multi-variant PDRA
@@ 5.1.2 Encoder
The encoder converts embedded input features into high-level contextual representations through a stack of $l$ transformer layers. In the traditional transformer encoder (Vaswani et al. 2017, Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024), each layer consists of two sublayers: (1) a multi-head attention (MHA) mechanism that models pairwise interactions between nodes, and (2) a position-wise feed-forward network (FFN) that independently refines the representation of each node. Each sublayer is followed by residual connections and instance normalization to ensure stable training and improved generalization. However, traditional transformer architectures face computational inefficiencies when processing large-scale routing problems, particularly in terms of memory usage and training stability. The standard attention mechanism exhibits quadratic complexity with sequence length, while conventional normalization techniques may not provide optimal gradient flow for deep networks. Additionally, traditional ReLU-based FFN have limited expressiveness for capturing complex non-linear relationships in routing optimization. To address these issues, our model adopts a modern transformer encoder (Berto et al. 2024), inspired by advancements in LLM (Nguyen et al. 2024, Dubey et al. 2024), with four key enhancements (illustrated in Figure 3):
@@
    Root Mean Square (RMS) normalization replaces instance normalization, providing better training stability and reducing computational overhead (Zhang and Sennrich 2019).
@@
    Pre-normalization configuration shifts the normalization step to before residual connections, enhancing gradient flow and accelerating convergence (Jiang et al. 2023).
@@
    FlashAttention (FA) is used in all MHA layers to improve efficiency when processing large input graphs without losing numerical accuracy (Dao et al. 2022, Dao 2023).
@@
    Swish Gated Linear Unit Feed-Forward Network (SGLUFFN), an extension of the Gated Linear Unit (Dauphin et al. 2017), replaces the traditional ReLU-based FFN, strengthening the model’s ability to capture complex non-linear relationships (Shazeer 2020).
@@
where $\mathbf{h}^{[l]}\in\mathbb{R}^{d}$ denotes the input vector to the $l$-th layer. Detailed formulations of RMS, FA and SGLUFFN are in Appendix A. Model component details, and ablation studies in Appendix Equation (B. Ablation study) validate the superiority of this modern architecture over traditional transformers.
@@ 5.1.3 Decoder
where $C$ acts as a clipping parameter for the tanh function, boosting solution exploration in the search space (Bello et al. 2016, Berto et al. 2024). Once the probability distribution $p_{i}$ is obtained, the current time step $t$ is concluded, and the state is updated to $t+1$, enabling the decoder to build solutions in an autoregressive manner that adapts to diverse PDRA variants. Detailed mathematical formulations of SHA and Softmax are also provided in Appendix A. Model component details.
@@ 5.2 Multi-task training
UM employs policy optimization with multiple optima (POMO) adapted for multi-task learning to simultaneously handle all PDRA variants within a single model (Kwon et al. 2020, Liu et al. 2024, Zhou et al. 2024, Berto et al. 2024). This approach eliminates the computational overhead of training separate models for each variant while leveraging shared knowledge across problem structures.
We implement a batch-wise strategy where each training batch focuses on consistent attribute configurations, ensuring uniform decoder sequence lengths within batches to facilitate efficient parallel processing and stable gradient computation (Liu et al. 2024, Zhou et al. 2024). Attributes are stochastically activated with 50% probability to enhance model robustness across diverse constraint configurations, allowing the model to learn adaptive strategies for both OR and TW scenarios. A critical challenge in multi-task learning is reward discrepancy across attribute combinations; for example, scenarios with TW may yield different reward scales compared to those without. To mitigate training biases, we implement a normalized reward strategy combining exponentially moving average (EMA) with Z-score normalization, following the approach established in Gong et al. (2025). This normalization anchors rewards to their historical trajectories within each attribute regime, enabling stable multi-task convergence across diverse PDRA variants. Algorithm 1 summarizes the complete multi-task training framework. The unified architecture learns shared representations across PDRA variants while adapting to variant-specific constraints through the attention mechanism and masking strategies, achieving superior performance compared to specialized single-task models.
@@ 6 Finetuning to unseen attributes
The finetuning mechanism leverages efficient adapter layers, drawing inspiration from recent advancements in parameter-efficient transfer learning (Berto et al. 2024). The core principle involves parameter augmentation with zero-initialized entries; this design choice ensures the preservation of pre-trained knowledge while accommodating new attribute-specific requirements. For integrating a new attribute, the finetuning mechanism requires only minimal modifications to the input embedding and the decoder’s context embedding. Taking MD as an example, these modifications are as follows. The original depot node embedding formulation:
@@ 7 Experiments
This section evaluates the proposed unified model through comprehensive experiments across multiple PDRA variants and network scales, examining solution quality, computational efficiency, scalability, and adaptability to unseen attributes.
@@ 7.1 Experiment setup and benchmarks
The experimental configuration encompasses model architecture specifications, problem parameter settings, training infrastructure, and benchmark methods.
Model Parameters. UM employs an embedding size $d=128$ and features a modern transformer encoder consisting of $l=6$ layers with 8-head attention and SGLUFFN with a hidden size of 512. These architectural settings are informed by successful applications in neural combinatorial optimization (Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Zhou et al. 2024, Liu et al. 2024, Berto et al. 2024, Gong et al. 2025). UM configuration yields approximately 1.3 million trainable parameters. Training spans 200 epochs (batch size = 64) on 100,000 on-the-fly generated instances to ensure robust learning across diverse problem variants. The Adam optimizer employs an initial learning rate of $10^{-4}$, $L_{2}$ regularization ($10^{-6}$ weight decay), and learning rate decay (×0.1 at epochs 175 and 195, respectively).
Problem Parameters. UM is trained on 100-node instances, comprising 50 nodes in set $P$ (with damage information values, representing 50 assessable road links) and 50 nodes in set $N$ (without damage information, including one depot). To evaluate generalization capability, testing is conducted on larger instances (200, 400, 600, 800, and 1,000 nodes), each with 50% nodes in $P$ and 50% in $N$. Damage information values for $P$ are uniformly sampled from [1, 10] and normalized by division by 10. Practically, drones have a 2-hour battery flight time limit $Q$ and a speed of 60 km/h (Zhang et al. 2023). Given the time-critical nature of disaster response, assessment time limits $p_{\max}$ are set to 30, 45, and 60 minutes (corresponding to 30, 45, and 60 km flight distances, respectively). Since $Q$ is significantly larger than $p_{\max}$, $Q$ is not considered in subsequent experiments, with a focus on $p_{\max}$. Road network coordinates are normalized to $[0,1]$, with link lengths scaled accordingly. For simplified data generation, training uses $p_{\max}=2,3,4$ corresponding to 30, 45, and 60 minutes in real-world distances. For parameter $K$, training uses 2, 3, or 4 drones. For the latest time windows parameter $l_{i}$ in the TW attribute, the generation method adopted is consistent with the approach proposed for classical vehicle routing problems in the existing literature (Zhao et al. 2020, Liu et al. 2024, Zhou et al. 2024, Berto et al. 2024).
Training Infrastructure. Training is conducted on Google Colab (https://colab.research.google.com) using an NVIDIA A100 GPU. Each epoch requires approximately 7.2 minutes, resulting in a total training duration of around 24 hours for 200 epochs.
Benchmarks. For comparison, two benchmarks are established to evaluate the approach as listed below:
@@
    Traditional Optimization Baseline: The mathematical model corresponding to each PDRA variant is detailed in Appendix C. Mathematical formulation, and each variant can thus be solved using commercial optimization solvers (e.g., Gurobi). Like UM, this approach requires no domain knowledge for algorithm design but necessitates separate model formulations and solving procedures for each of the 8 variants, highlighting the computational overhead UM aims to eliminate. Gurobi is evaluated with multiple time limits (60s, 10×60s, 20×60s, 30×60s) to reflect different urgency levels in disaster response scenarios.
@@
    Single-Task Deep Learning Baseline: Comparisons are made with single-task DRL methods from the literature, specifically the AEDM proposed by Gong et al. (2025). These include:
    • AEDM-Basic: Single-task model for PDRA-Basic
    • AEDM-OR: Single-task model for PDRA-OR
    • AEDM-TW: Single-task model for PDRA-TW
    • AEDM-OR-TW: Single-task model for PDRA-OR-TW
    • AEDM-MD: Single-task model for PDRA-MD
    • AEDM-OR-MD: Single-task model for PDRA-OR-MD
    • AEDM-TW-MD: Single-task model for PDRA-TW-MD
    • AEDM-OR-TW-MD: Single-task model for PDRA-OR-TW-MD.
Each baseline requires separate training for specific problem variants, necessitating multiple models for comprehensive coverage. In contrast, UM handles all 8 variants within a single model. As shown in Table 3, the existing literature approach requires training 8 separate models (one for each variant) with a total of 10.4 million parameters across 1,600 epochs, resulting in 192 hours of training time. UM achieves comprehensive coverage with a single 1.3 million parameter model trained for 200 epochs in just 24 hours, representing an 8× reduction in model count, parameter count, training epochs, and training time.
@@
Since post-disaster response requires assessing road networks at multiple affected locations (e.g., after earthquakes or floods), the task is inherently multi-instance rather than single-instance. Therefore, Gurobi, AEDM, and the proposed UM must all be evaluated on their ability to handle multiple instances efficiently, making both solution quality and computational speed equally critical. For this reason, the following results are averaged over 10 randomly generated problem instances. For the single-task AEDM and UM, we report the total computation time across the 10 instances, while for Gurobi we report the average time per instance, which places Gurobi in a relatively favorable position.
Table 3: Comparison of Methodological Efficiencies
Method No. of models Total Parameters (M) Total Epochs Training Cost (h)
Existing literature 8 10.4 1,600 192
Ours 1 1.3 200 24
@@ 7.2 Computational performance
AEDM-Basic 2 18.56 4.43% 2 25.87 5.45%
AEDM-OR 2 19.35 0.36% 2 27.28 0.29%
AEDM-TW 2 17.02 12.36% 2 23.04 15.79%
AEDM-OR-TW 2 17.68 8.96% 2 24.34 11.04%
UM 2 19.42 0.00% 2 27.36 0.00%
@@
AEDM-Basic 2 13.59 5.10% 2 18.77 3.40%
AEDM-OR 2 12.40 13.41% 2 16.98 12.61%
AEDM-TW 2 14.40 -0.56% 2 19.34 0.46%
AEDM-OR-TW 2 14.22 0.70% 2 18.47 4.94%
UM 2 14.32 0.00% 2 19.43 0.00%
@@
AEDM-Basic 2 15.11 11.01% 2 20.86 11.20%
AEDM-OR 2 15.74 7.30% 2 21.36 9.07%
AEDM-TW 2 15.51 8.66% 2 21.20 9.75%
AEDM-OR-TW 2 16.76 1.30% 2 23.10 1.66%
UM 2 16.98 0.00% 2 23.49 0.00%
@@
AEDM-Basic 1 64.16 -0.66% 1 72.42 0.47%
AEDM-OR 1 58.21 8.68% 1 65.28 10.28%
AEDM-TW 1 44.35 30.42% 1 47.23 35.09%
AEDM-OR-TW 1 42.47 33.37% 1 46.43 36.19%
UM 1 63.74 0.00% 1 72.76 0.00%
@@
10×60 29.19 59.89% 10×60 37.17 54.87%
20×60 48.40 33.49% 20×60 44.23 46.30%
30×60 52.97 27.21% 30×60 51.94 36.94%
AEDM-Basic 1 60.97 16.22% 1 74.83 9.14%
AEDM-OR 1 72.55 0.30% 1 82.27 0.11%
AEDM-TW 1 45.01 38.15% 1 47.97 41.76%
AEDM-OR-TW 1 52.28 28.16% 1 50.40 38.81%
UM 1 72.77 0.00% 1 82.36 0.00%
@@
AEDM-Basic 1 39.90 8.63% 1 46.07 8.54%
AEDM-OR 1 34.14 21.82% 1 40.47 19.65%
AEDM-TW 1 42.16 3.46% 1 49.80 1.13%
AEDM-OR-TW 1 41.38 5.24% 1 48.37 3.97%
UM 1 43.67 0.00% 1 50.37 0.00%
@@
AEDM-Basic 1 40.37 8.48% 1 46.98 7.56%
AEDM-OR 1 39.55 10.34% 1 46.46 8.58%
AEDM-TW 1 42.00 4.78% 1 48.84 3.90%
AEDM-OR-TW 1 43.33 1.77% 1 49.00 3.58%
UM 1 44.11 0.00% 1 50.82 0.00%
Tables 4 and 5 present the numerical experiments of the comparisons for 200-node and 400-node networks, where the column of gap is calculated as $\text{Gap}=(y-y_{\text{other}})/y$, where $y$ denotes the objective value of UM and $y_{\text{other}}$ that of the comparison method. From these results, two key insights emerge:
@@
    Solution Quality Superiority. UM consistently achieves the highest objective values across all PDRA variants. For the traditional commercial solver (Gurobi), solution quality gradually improves as the time limit extends; however, even with the longest time limit (30×60s, 30 minutes), its final solution quality remains suboptimal compared to UM. In 400-node networks, Gurobi frequently fails to generate feasible solutions within practical time limits (marked by ✗ in Table 5). For single-task AEDMs, high solution quality is only achieved when the model is specifically trained for the target PDRA variant. If a single-task AEDM is applied to a non-matching variant (e.g., using AEDM-Basic for PDRA-OR-TW), its solution quality drops significantly.
@@
    Computational Efficiency. UM delivers real-time performance, generating valid solutions within 1–2 seconds across all tested network scales and drone fleet sizes. In contrast, Gurobi requires 10–30 minutes to reach reasonable-quality solutions, far exceeding the time-critical demands of post-disaster response. Single-task AEDMs also suffer from inefficiency. They require separate models (one for each PDRA variant), leading to an increase in total training time and necessitating variant-specific deployment, which further complicates real-world application.
Figures 4 and 5 visually corroborate these findings. The radar charts in Figure 4 show UM consistently occupying the outermost (best) performance positions across all parameter settings and problem variants, indicating superior and stable performance under diverse operational constraints. Figure 5 shows that UM consistently outperforms baseline methods in terms of solution quality (total collected information value). Specifically, UM surpasses the best single-task model by 6%–14% and Gurobi by 24%–82% across tested configurations.
Figure 4: Average Performance Gap Calculation
Figure 5: UM vs. Baseline Methods: Average Gap Computation

### v2

@@ 5.1 Model architecture
Figure 2: Architecture of the Unified Model for Multi-Task PDRA. The unified model processes road network data, problem parameters, and variant attributes through five main components: (1) Input layer embeds node features and problem configurations, distinguishing road network nodes and depot nodes; (2) Encoder transforms embeddings into high-level contextual representations using modern transformer layers with RMS normalization, FlashAttention, and SGLUFFN; (3) Decoder constructs solutions autoregressively through probability computation, softmax selection, and single-head attention mechanisms; (4) Update module maintains solution feasibility by managing infeasible action masks, active drone indices, and current time states; (5) Output produces feasible multi-drone routes satisfying variant-specific constraints with appropriate termination conditions.
@@ 5.1.2 Encoder
The encoder converts embedded input features into high-level contextual representations through a stack of $l$ transformer layers. In the traditional transformer encoder (waswani2017attention; kool2018attention; kwon2020pomo; zhang2020multi; luo2023neural; zhou2024mvmoe; liu2024multi), each layer consists of two sublayers: (1) a multi-head attention (MHA) mechanism that models pairwise interactions between nodes, and (2) a position-wise feed-forward network (FFN) that independently refines the representation of each node. Each sublayer is followed by residual connections and instance normalization to ensure stable training and improved generalization. However, traditional transformer architectures face computational inefficiencies when processing large-scale routing problems, particularly in terms of memory usage and training stability. The standard attention mechanism exhibits quadratic complexity with sequence length, while conventional normalization techniques may not provide optimal gradient flow for deep networks. Additionally, traditional ReLU-based FFN have limited expressiveness for capturing complex non-linear relationships in routing optimization. To address these issues, our model adopts a modern transformer encoder (berto2024routefinder), inspired by advancements in LLM (nguyen2024sequence; dubey2024llama), with four key enhancements (illustrated in Figure 3):
@@
    Root Mean Square (RMS) normalization replaces instance normalization, providing better training stability and reducing computational overhead (zhang2019root).
@@
    Pre-normalization configuration shifts the normalization step to before residual connections, enhancing gradient flow and accelerating convergence (jiang2023pre).
@@
    FlashAttention (FA) is used in all MHA layers to improve efficiency when processing large input graphs without losing numerical accuracy (dao2022flashattention; dao2023flashattention).
@@
    Swish Gated Linear Unit Feed-Forward Network (SGLUFFN), an extension of the Gated Linear Unit (dauphin2017language), replaces the traditional ReLU-based FFN, strengthening the model’s ability to capture complex non-linear relationships (shazeer2020glu).
@@
where $\mathbf{h}^{[l]}\in\mathbb{R}^{d}$ denotes the input vector to the $l$-th layer. Detailed formulations of RMS, FA and SGLUFFN are in Appendix A, and ablation studies in Appendix Equation (B) validate the superiority of this modern architecture over traditional transformers.
@@ 5.1.3 Decoder
where $C$ acts as a clipping parameter for the tanh function, boosting solution exploration in the search space (bello2016neural; berto2024routefinder). Once the probability distribution $p_{i}$ is obtained, the current time step $t$ is concluded, and the state is updated to $t+1$, enabling the decoder to build solutions in an autoregressive manner that adapts to diverse PDRA variants. Detailed mathematical formulations of SHA and Softmax are also provided in Appendix A.
@@ 5.2 Multi-task training
UM employs policy optimization with multiple optima (POMO) adapted for multi-task learning to simultaneously handle all PDRA variants within a single model (kwon2020pomo; liu2024multi; zhou2024mvmoe; berto2024routefinder). This approach eliminates the computational overhead of training separate models for each variant while leveraging shared knowledge across problem structures.
We implement a batch-wise strategy where each training batch focuses on consistent attribute configurations, ensuring uniform decoder sequence lengths within batches to facilitate efficient parallel processing and stable gradient computation (liu2024multi; zhou2024mvmoe). Attributes are stochastically activated with 50% probability to enhance model robustness across diverse constraint configurations, allowing the model to learn adaptive strategies for both OR and TW scenarios. This stochastic activation balances training exposure and prevents bias toward any specific attribute. A critical challenge in multi-task learning is reward discrepancy across attribute combinations; for example, scenarios with TW may yield different reward scales compared to those without. To mitigate training biases, we implement a normalized reward strategy combining exponentially moving average (EMA) with Z-score normalization, following the approach established in Gong2025Deep. This normalization anchors rewards to their historical trajectories within each attribute regime, enabling stable multi-task convergence across diverse PDRA variants. Algorithm 1 summarizes the complete multi-task training framework. The unified architecture learns shared representations across PDRA variants while adapting to variant-specific constraints through the attention mechanism and masking strategies, achieving superior performance compared to specialized single-task models.
@@ 6 Finetuning to unseen attributes
The finetuning mechanism leverages efficient adapter layers, drawing inspiration from recent advancements in parameter-efficient transfer learning (berto2024routefinder). The core principle involves parameter augmentation with zero-initialized entries; this design choice ensures the preservation of pre-trained knowledge while accommodating new attribute-specific requirements. For integrating a new attribute, the finetuning mechanism requires only minimal modifications to the input embedding and the decoder’s context embedding. Taking MD as an example, these modifications are as follows. The original depot node embedding formulation:
@@ 7 Experiments
This section evaluates the proposed UM through extensive experiments across multiple PDRA variants and network scales, examining solution quality, computational efficiency, scalability, and adaptability to unseen attributes.
@@ 7.1 Experiment setup and benchmarks
The experimental configuration encompasses model architecture specifications, problem parameter settings, training infrastructure, benchmark methods, and evaluation protocol.
Model Parameters. UM employs an embedding size $d=128$ and features a modern transformer encoder consisting of $l=6$ layers with 8-head attention and SGLUFFN with a hidden size of 512. These architectural settings are informed by successful applications in neural combinatorial optimization (kool2018attention; kwon2020pomo; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder; Gong2025Deep). UM configuration yields approximately 1.3 million trainable parameters. Training spans 200 epochs (batch size = 64) on 100,000 on-the-fly generated instances to ensure robust learning across diverse problem variants. The Adam optimizer employs an initial learning rate of $10^{-4}$, $L_{2}$ regularization ($10^{-6}$ weight decay), and learning rate decay (×0.1 at epochs 175 and 195, respectively).
Problem Parameters. UM is trained on 100-node instances, comprising 50 nodes in set $P$ (with damage information values, representing 50 assessable road links) and 50 nodes in set $N$ (without damage information, including one depot). To evaluate generalization capability, testing is conducted on larger instances (200, 400, 600, 800, and 1,000 nodes), each with 50% nodes in $P$ and 50% in $N$. Damage information values for $P$ are uniformly sampled from [1, 10] and normalized by division by 10. Practically, drones have a 2-hour battery flight time limit $Q$ and a speed of 60 km/h (zhang2023robust). Given the time-critical nature of disaster response, assessment time limits $p_{\max}$ are set to 30, 45, and 60 minutes (corresponding to 30, 45, and 60 km flight distances, respectively). Since $Q$ is significantly larger than $p_{\max}$, $Q$ is not considered in subsequent experiments, with a focus on $p_{\max}$. Road network coordinates are normalized to $[0,1]$, with link lengths scaled accordingly. For simplified data generation, training uses $p_{\max}=2,3,4$ corresponding to 30, 45, and 60 minutes in real-world distances. For parameter $K$, training uses 2, 3, or 4 drones. For the latest time windows parameter $l_{i}$ in the TW attribute, the generation method adopted is consistent with the approach proposed for classical vehicle routing problems in the existing literature (zhao2020hybrid; liu2024multi; zhou2024mvmoe; berto2024routefinder).
Training Infrastructure. Training is conducted on Google Colab (https://colab.research.google.com) using an NVIDIA A100 GPU. Each epoch requires approximately 7.2 minutes, resulting in a total training duration of around 24 hours for 200 epochs. The source code for UM is publicly available at https://github.com/PJ-HTU/UM_PDRA.
Benchmark Methods. To extensively evaluate UM’s performance, we establish three categories of baseline methods representing different optimization paradigms: exact optimization, heuristic optimization, and single-task deep learning.
@@
    Exact Optimization Baseline. The mathematical formulation for each PDRA variant is provided in Appendix C. We solve these formulations using Gurobi, a state-of-the-art commercial optimizer. While the exact method requires no algorithm design expertise, it necessitate variant-specific model formulations due to differing constraint structures across OR, TW, and MD attributes. To reflect varying urgency levels in disaster response scenarios, Gurobi is evaluated under two time limit settings: 60 seconds (Gurobi-1×60) and 1,800 seconds (Gurobi-30×60), with the best solution reported for each setting.
@@
    Heuristic Optimization Baseline. We develop a family of two-phase heuristics specifically designed for PDRA variants, with detailed algorithmic descriptions provided in Appendix D. These heuristics are inspired by classical OP frameworks (kobeaga2018efficient; yang2025heuragenix) but are substantially modified to accommodate the unique structural and feasibility characteristics of PDRA. Each heuristic consists of: (1) a greedy construction (GC) phase that builds initial solutions using a composite scoring function that balances information value collection with temporal urgency, adapted to the transformed network structure where artificial nodes represent road links with exclusive visit requirements, and (2) a local search (LS) improvement phase that applies three operators (relocate, exchange, and remove-insert) carefully adapted to handle artificial-node triplet structures while respecting variant-specific constraints. Due to fundamental differences in feasibility rules across variants (OR modifies depot return logic, TW enforces temporal constraints, and MD introduces depot assignment requirements), each variant requires its own tailored heuristic implementation. We denote these heuristics as GC-Variant for construction only and GC+LS-Variant for construction with improvement, where Variant $\in$ {Basic, OR, TW, OR-TW, MD, OR-MD, TW-MD, OR-TW-MD}, corresponding to all eight PDRA variants. While these heuristics provide strong baselines, their requirement for hand-crafted domain knowledge and variant-specific redesign exemplifies the limitations that motivate UM’s unified learning approach.
  3. c)
    Single-Task Deep Learning Baseline. We compare UM against single-task attention-based encoder-decoder models (AEDM) from Gong2025Deep. Following the single-task paradigm, eight independent AEDM models are trained, one for each PDRA variant: AEDM-Basic, AEDM-OR, AEDM-TW, AEDM-OR-TW, AEDM-MD, AEDM-OR-MD, AEDM-TW-MD, and AEDM-OR-TW-MD. Each model undergoes separate training (200 epochs per variant) to learn variant-specific routing strategies. This single-task approach results in high training costs and deployment complexity: 8 separate models with a total of 10.4 million parameters, 1,600 cumulative training epochs, and 192 hours of training time.
As summarized in Table 3, the three baseline categories demonstrate distinct limitations that motivate our unified framework. Exact methods require variant-specific mathematical formulations for each PDRA configuration, resulting in high deployment complexity and poor computational scalability despite minimal algorithmic expertise requirements. Heuristic methods demand extensive domain knowledge and variant-specific algorithmic redesign, with each of the eight PDRA variants necessitating tailored implementations that complicate practical deployment. Single-task DRL approaches eliminate the need for domain expertise and achieve rapid performance, but require training separate models for each variant, leading to substantial deployment overhead when comprehensive variant coverage is needed. In contrast, UM consolidates all 8 variants within a single 1.3M parameter model trained for 200 epochs in 24 hours, achieving an 8-fold reduction in training cost while enabling streamlined deployment through unified problem representation.
@@
Evaluation Protocol. Since post-disaster response inherently involves assessing road networks at multiple affected locations (e.g., after earthquakes or floods), we evaluate all methods on their ability to handle multiple problem instances efficiently, where both solution quality and computational speed are critical. Results are averaged over 10 randomly generated instances for each problem configuration. For UM and AEDM, we report total computation time across all 10 instances to reflect practical deployment scenarios, while for Gurobi and heuristics we report average time per instance, which provides these baseline methods a favorable computational comparison basis.
Table 3: Comparison of Methodological Requirements and Efficiency
Method
Models/Formulations
per Variant
Deployment
Complexity
Total Training
(for 8 variants)
Domain
Expertise
Exact optimization baseline 8 formulations High N/A Low
Heuristic optimization baseline 8 heuristics High N/A High
Single-task deep learning baseline 8 models High 10.4M params, 192h Low
UM (Ours) 1 model Low 1.3M params, 24h Low
@@ 7.2 Computational performance
GC-OR 1 14.41 25.78% 2 16.38 40.13%
GC+LS-OR 17 16.31 16.01% 39 20.14 26.39%
AEDM-Basic 1 18.56 4.43% 1 25.87 5.45%
AEDM-OR 1 19.35 0.36% 1 27.28 0.29%
AEDM-TW 1 17.02 12.36% 1 23.04 15.79%
AEDM-OR-TW 1 17.68 8.96% 1 24.34 11.04%
UM 1 19.42 0.00% 1 27.36 0.00%
@@
GC-TW 2 10.51 26.61% 5 14.48 25.48%
GC+LS-TW 23 13.14 8.24% 51 17.41 10.40%
AEDM-Basic 1 13.59 5.10% 1 18.77 3.40%
AEDM-OR 1 12.40 13.41% 1 16.98 12.61%
AEDM-TW 1 14.40 -0.56% 1 19.34 0.46%
AEDM-OR-TW 1 14.22 0.70% 1 18.47 4.94%
UM 1 14.32 0.00% 1 19.43 0.00%
@@
GC-OR-TW 2 11.34 33.22% 7 12.56 46.53%
GC+LS-OR-TW 24 13.94 17.91% 58 16.85 28.27%
AEDM-Basic 1 15.11 11.01% 1 20.86 11.20%
AEDM-OR 1 15.74 7.30% 1 21.36 9.07%
AEDM-TW 1 15.51 8.66% 1 21.20 9.75%
AEDM-OR-TW 1 16.76 1.30% 1 23.10 1.66%
UM 1 16.98 0.00% 1 23.49 0.00%
@@
GC-Basic 15 40.21 36.90% 30 42.60 41.46%
GC+LS-Basic 350 51.21 19.65% 856 53.12 27.00%
AEDM-Basic 2 64.16 -0.66% 2 72.42 0.47%
AEDM-OR 2 58.21 8.68% 2 65.28 10.28%
AEDM-TW 2 44.35 30.42% 2 47.23 35.09%
AEDM-OR-TW 2 42.47 33.37% 2 46.43 36.19%
UM 2 63.74 0.00% 2 72.76 0.00%
@@
30×60 52.97 27.21% 30×60 37.17 54.87%
GC-OR 17 46.40 36.23% 28 31.42 61.85%
GC+LS-OR 361 55.14 24.23% 841 60.69 26.31%
AEDM-Basic 2 60.97 16.22% 2 74.83 9.14%
AEDM-OR 2 72.55 0.30% 2 82.27 0.11%
AEDM-TW 2 45.01 38.15% 2 47.97 41.76%
AEDM-OR-TW 2 52.28 28.16% 2 50.40 38.81%
UM 2 72.77 0.00% 2 82.36 0.00%
@@
GC-TW 23 15.11 65.40% 31 21.21 57.89%
GC+LS-TW 469 30.16 30.94% 958 36.24 28.06%
AEDM-Basic 2 39.90 8.63% 2 46.07 8.54%
AEDM-OR 2 34.14 21.82% 2 40.47 19.65%
AEDM-TW 2 42.16 3.46% 2 49.80 1.13%
AEDM-OR-TW 2 41.38 5.24% 2 48.37 3.97%
UM 2 43.67 0.00% 2 50.37 0.00%
@@
GC-OR-TW 24 15.20 65.55% 35 14.50 71.47%
GC+LS-OR-TW 620 23.12 47.60% 1104 22.18 56.36%
AEDM-Basic 2 40.37 8.48% 2 46.98 7.56%
AEDM-OR 2 39.55 10.34% 2 46.46 8.58%
AEDM-TW 2 42.00 4.78% 2 48.84 3.90%
AEDM-OR-TW 2 43.33 1.77% 2 49.00 3.58%
UM 2 44.11 0.00% 2 50.82 0.00%
Figure 4: Comparative Performance of Methods Across Problem Variants
Figure 5: Performance Gap and Solution Time Distribution Across Methods
Tables 4 and 5 present the numerical experiments of the comparisons for 200-node and 400-node networks, where the relative performance gap is calculated as $\text{Gap}=(y-y_{\text{other}})/y$, where $y$ denotes the objective value of UM and $y_{\text{other}}$ that of the comparison method. From these results, two key insights emerge:
@@
    Solution Quality Superiority. UM consistently achieves the highest objective values across all PDRA variants. For the commercial solver (Gurobi), solution quality gradually improves as the time limit extends; however, even with the longest time limit (1,800 seconds), its final solution quality remains suboptimal compared to UM. In 400-node networks, Gurobi frequently fails to generate feasible solutions within practical time limits (marked by ✗ in Table 5). For heuristic methods, GC+LS achieves better performance than GC alone through local search improvement, but both remain substantially inferior to UM. For single-task AEDMs, high solution quality is only achieved when the model is specifically trained for the target PDRA variant; when applied to non-matching variants (e.g., using AEDM-Basic for PDRA-OR-TW, which is feasible due to identical model architectures), performance degrades significantly.
@@
    Computational Efficiency. UM delivers rapid performance, generating valid solutions within 1–2 seconds across all tested network scales and drone fleet sizes. In contrast, Gurobi requires 1,800 seconds, (30 minutes) to reach reasonable-quality solutions, far exceeding the time-critical demands of post-disaster response. Heuristic methods exhibit varying computational times: GC provides rapid construction but suboptimal quality, while GC+LS trades additional computational time (ranging from seconds to minutes depending on problem complexity) for improved solutions that still fall short of UM’s performance. Single-task AEDMs also suffer from inefficiency as they require separate models (one for each PDRA variant), leading to increased total training time and necessitating variant-specific deployment, which further complicates real-world application.
Figures 4 and 5 visually corroborate these findings. The radar charts in Figure 4 show UM consistently occupying the outermost (best) performance positions across all parameter settings and problem variants, indicating superior and stable performance under diverse operational constraints. Figure 5 (a) shows that UM consistently outperforms baseline methods in terms of solution quality (total collected information value). Specifically, UM surpasses the best single-task DRL models by 6%–14%, heuristic methods (GC+LS) by 22%–42%, and Gurobi by 24%–82% across tested configurations. Figure 5 (b) demonstrates the computational efficiency advantages: UM maintains rapid performance with solution times of 1–2 seconds, comparable to single-task AEDM, while substantially outperforming heuristic methods (12–274 seconds) and Gurobi (60–1,800 seconds).